### v3

# ESG Rating Disagreement and Corporate Total Factor Productivity:Inference and Prediction

Abstract
This paper examines how ESG rating disagreement ($Dis$) affects corporate total factor productivity (TFP) in China based on data of A-share listed companies from 2015 to 2022. We find that $Dis$ reduces TFP, especially in state-owned, non-capital-intensive, and low-pollution firms. Mechanism analysis shows that green innovation strengthens the dampening effect of $Dis$ on TFP, and that $Dis$ lowers corporate TFP by increasing financing constraints. Furthermore, XGBoost regression demonstrates that $Dis$ plays a significant role in predicting TFP, with SHAP showing that the dampening effect of ESG rating disagreement on TFP is still pronounced in firms with large $Dis$ values.
JEL Codes: G32, G24, C54.
Keywords: ESG rating disagreement; Total factor productivity; Green innovation; Financing constraints; Machine learning
1 Introduction
Since 1987, the United Nations has reinforced its commitment to sustainable development, and introduced the concept of ESG (Environmental, Social, and Governance) by then-UN Secretary-General Kofi Annan in 2004. Since then, ESG principles have been increasingly integrated into global investment strategies and corporate assessments, with their specific meanings evolving over time. In recent years, China is prominently advocating the philosophy that "green mountains and clear waters are as valuable as mountains of gold and silver", and promoting sustainable development both domestically and internationally. On December 27, 2023, the Guidelines for the Comprehensive Advancement of Beautiful China Construction was issued by the Central Committee of the Communist Party of China and the State Council clearly clarified the exploration and development of ESG assessment frameworks, which shows a strong will of establishing ESG Evaluation System with Chinese Characteristics.
Despite the implementation of various ESG policies in recent year, there are still significant disagreement in ESG rating system. From the moment the ESG concept was introduced in China, multiple rating agents have independently published their ESG rating reports. However, these reports often display significant discrepancies in ratings, which undoubtedly has a substantial impact on the companies being evaluated. Examining the economic consequences of these ESG rating disagreements is of considerable value for both enterprises and regulatory bodies. The main recent studies are: Wang et al. (2024); Zhao et al. (2024) argue that disagreements in ESG ratings are significantly negatively correlated with stock returns; ESG rating disagreement increase analyst forecast errors(Liu et al., 2024); ESG rating disagreement drives firms towards green innovation (Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); Moreover, increased disagreement in ESG ratings leads to higher levels of real earnings management both in the current period and over the next one to two years (Li and Cheng, 2024); The firms with agreed high ESG rating can benefit through cost reduction and larger amount debt-financing(Guo et al., 2024); Besides,the disagreement in ESG ratings may poses a potential risk of stock price collapse(Luo et al., 2023); What’s more, ESG rating disagreement significantly increases audit fees due to information asymmetry(Ling et al., 2024).
However, the focus of most recent studies has been less on exploring the impact of ESG rating disagreements on the overall development of a company. Therefore, we select TFP as the subject of our analysis in this study, as it serves as an indicator of a company’s high-quality development.
2 Theoretical Analysis and Research Hypotheses
According to information asymmetry theory, ESG rating disagreements can confuse stakeholders and contribute to the degree of information asymmetry between companies and investors. ESG ratings, as non-financial information, provide stakeholders with a more comprehensive insight into companies, thereby alleviating information asymmetry(Xiaohong and Zhenghan, 2023). Conversely, significant ESG rating disagreements can intensify information asymmetry, creating a noise effect that leads to investor misjudgments(Berg et al., 2022). Thereby, increasing financing constraints and ultimately resulting in a decline in total factor productivity. Simultaneously, based on stakeholder theory, companies may take certain remedial measures in response to ESG rating disagreements to balance the interests of various stakeholders. For example, companies might intensify green innovation under the stimulus of ESG rating disagreements(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); However, as this initiative cannot cater to all stakeholders under noise effect caused by ESG rating disagreement, it may lead some groups to perceive "greenwashing" behaviors, which ultimately undermines the company’s financing efforts and results in a decline in total factor productivity(Peng and Xie, 2024).
In summary, we propose the research hypothesis H1: ESG rating disagreement reduces corporate TFP.
3 Data and Empirical Design
3.1 Sample Selection
We utilizes financial data of A-share listed companies in China from 2015 to 2022111Before 2015, the only more well-known local ESG rating agencies in China were Huazheng, alongside ESG ratings from four agencies: Huazheng, Wind, SynTao Green Finance, and MioTech. The data processing procedures follow these steps: (1) excluding companies in the financial and real estate industries; (2) excluding ST, *ST, and delisted companies; (3) excluding companies with only one ESG rating in a given year; (4) excluding companies with missing key indicators; and (5) applying Winsorization at the 1% level on both tails. All data is sourced from the CSMAR database, the CNRDS database, and the four rating agencies, resulting in 13417 samples in baseline regression.
3.2 Model Specification
To test the hypotheses presented earlier,we conduct the following baseline regression model [^2] Considering the time required for information to be transmitted and for the market to react, TFP is lagged by one period:
$TFP_{i,t+1}=\alpha_{0}+\alpha_{1}Dis_{i,t}+\alpha_{2}Control_{i,t}+Year_{t}+Id_{i}+\epsilon_{i,t}$ (1)
where $TFP_{i,t+1}$ is the total factor productivity of company $i$ in year $t+1$, $Dis_{i,t}$ is the ESG rating discrepancy, $Control_{i,t}$ represents the control variables, $Year_{t}$ and $Id_{i}$ denote year and company fixed effects, respectively, and $\epsilon_{i,t}$ is the error term.
3.3 Variable Descriptions
(1) Dependent Variable
The dependent variable in this study is Corporate Total Factor Productivity. Considering the robustness of the empirical process, we use three methods to measure TFP, they are LP introduced byLevinsohn and Petrin (2003) and adopted by Lu and Lian (2012), OP by Olley and Pakes (1992) and GMM by Blundell and Bond (1998). In the prediction task, we use TFP measured by LP method as the dependent variable(Xue et al., 2024).
(2) Independent Variable
The independent variable in this study is ESG Rating Disagreement. Domestic ESG rating agencies have advantages in terms of time and spatial proximity that foreign agencies cannot match, making them more likely to influence the behavior of domestic firms. Therefore, this study uses ESG rating data from four Chinese agencies: Huazheng, Wind, SynTao Green Finance, and MioTech. Following Avramov et al. (2022), the ratings were assigned positive values according to their rating levels, with a difference of 1 between levels. We define the rankings as the values normalized to the range [0, 1] for the given rating agency in a given year, and the pairwise standard deviation of these rankings is used as the measure of $Dis$.
$Dis_{i,t}=\sqrt{\frac{1}{n-1}\sum_{k\neq j}^{n}(rank_{i,t,k}-rank_{i,t,j})^{2}}$ (2)
where $rank_{i,j,k}$ represents the normalized ranking of corporate $i$ in year $t$ given by agency $k$, $n$ represents the number of pairs, $k$ and $j$ represent different ESG rating agencies.
(3) Control Variables
Based on the characteristics of our research and with reference to the works of Yang et al. (2024) and Xue et al. (2024), the control variables include firm value, top 10 shareholders’ ownership percentage, years listed, leverage ratio, firm size, revenue per capita, and ownership ratio. The main variables are defined in Table 1.
Table 1: Definition of Main Variables
Symbol Source Description
$TFP$ CSAMR Estimated using the LP, OP and GMM method
$Dis$ ESG Rating Agencies Calculated based on the method described above
$TobinQ$ CSAMR Tobin’s Q ratio
$Top10$ CSAMR Shareholding of top 10 shareholders / total shares
$ListAge$ CSAMR Natural logarithm of years since the firm’s IPO
$Lev$ CSAMR Total liabilities at year-end / total assets at year-end
$Size$ CSAMR Natural logarithm of total assets at year-end
$Avg$ CSAMR Natural logarithm of operating income / number of employees
$Der$ CSAMR Total liabilities at year-end / total equity at year-end
$EI$ CNRDS Natural logarithm of the number of green patents obtained in a given year
$FC$ CSAMR WW and KZ indices, the larger they are the more constrained the financing is
4 Empirical Analysis
Table 2: Descriptive Statistics of Main Variables
Variables Count Mean SD Min 25% 50% 75% Max
$TFP_{LP}$ 13417 8.51 1.04 6.38 7.79 8.40 9.14 11.28
$TFP_{OP}$ 13417 6.83 0.85 5.14 6.24 6.72 7.33 9.17
$TFP_{GMM}$ 13417 5.72 0.80 4.09 5.18 5.63 6.18 8.12
$Dis$ 13417 0.26 0.19 0.01 0.11 0.24 0.34 0.83
$TobinQ$ 13417 2.03 1.32 0.83 1.22 1.61 2.31 8.34
$Top10$ 13417 56.77 15.15 23.59 45.91 56.95 67.87 90.89
$ListAge$ 13417 2.34 0.68 1.10 1.79 2.40 3.00 3.37
$Lev$ 13417 0.43 0.19 0.06 0.28 0.43 0.57 0.88
$Size$ 13417 22.52 1.34 20.09 21.54 22.32 23.31 26.41
$Avg$ 13417 4.79 0.80 3.07 4.25 4.69 5.25 7.17
$Der$ 13417 1.05 1.07 0.07 0.39 0.74 1.31 6.75
$EI$ 13417 1.30 1.49 0.00 0.00 1.10 2.30 7.74
$WW$ 10059 -1.03 0.08 -1.24 -1.08 -1.03 -0.98 -0.86
$KZ$ 13417 1.17 1.94 -4.35 0.02 1.36 2.49 6.22
  • Note: The data in the table are descriptive statistics for deleting missing values after lagging one period of TFP.
4.1 Baseline Regression
Table 3 reports the results of the baseline regression, with clustering at the individual and time levels. The results indicate that ESG rating disagreement reduces corporate total factor productivity, providing preliminary support for the validity of Hypothesis H1.
Table 3: Baseline Regression Results
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.0675*** -0.0536** -0.0536** -0.0447*** -0.0357** -0.0373**
(-2.7452) (-2.1838) (-2.2098) (-2.9449) (-2.2503) (-2.2395)
$TobinQ$0.0032 -0.0017 -0.0007
(-0.5471) (-0.3953) (-0.1795)
$Top10$0.0010 -0.0009 -0.0005
(-0.9976) (-0.9496) (-0.4838)
$ListAge$ 0.0989** 0.0855** 0.0160
(2.4952) (2.3667) (0.4241)
$Lev$0.1697 -0.1557 -0.1481
(-1.5121) (-1.5361) (-1.4348)
$Size$ 0.3737*** 0.2581*** 0.2245***
(13.129) (14.214) (10.939)
$Avg$ 0.2292*** 0.2695*** 0.2671***
(2.7852) (2.8274) (2.7426)
$Der$$0*$0190* -0.0211**
(-2.0405) (-1.8342) (-1.9816)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
Doubel Clustering YES YES YES YES YES YES
$N$ 13417 13417 13417 13417 13417 13417
$R^{2}$ 0.0015 0.0011 0.0011 0.1785 0.1670 0.1409
  • Note: t-statistics are in parentheses. * p<0.1, ** p<0.05, *** p<0.01. The same applies below.
4.2 Robustness Tests
To ensure the robustness of the baseline regression results, the following robustness tests were conducted.
First, to verify whether Hypothesis H1 holds when only leading $Dis$ by one period, we lead the it by one period(see table8).
Second, since the total factor productivity of firms in this study is influenced by regional productivity levels, to mitigate endogeneity issues arising from omitted variable bias, we include the provincial-level New quality productivity measured by Lu et al. (2024) as a control variable (see table9).
Third, following Yang et al. (2022), we calculate the industry average of ESG rating disagreements by year and industry, and use the cube of the difference between the industry average and the ESG rating disagreements as an instrumental variable. A two-stage least squares (2SLS) estimation is then conducted to further mitigate potential endogeneity issues (see table10).
Fourth, we denote disagreements under 20% as non-existent and the remainder as present. Propensity scores are computed via logistic regression, followed by 1:1 nearest-neighbor matching and subsequent regression (see table11).
4.3 Mechanism Analysis
4.3.1 The Role of Green Innovation
Myriad existing researches demonstrate that ESG rating disagreement can drive firms to enhance green innovation(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024b). However, does this green innovation truly contribute to substantial corporate development? This study introduces green innovation as a mechanism variable, measured by the natural logarithm of the number of green patents held by a company. To examine the role of green innovation in the relationship between ESG rating disagreement and TFP, an interaction term between green innovation and ESG rating disagreement is constructed, with other settings remaining consistent with Equation 1.
Table 4 and figure1 reports the results. The results show that when there are no ESG rating discrepancies, green innovation promotes the improvement of TFP. Conversely, when there are ESG rating discrepancies, it strengthens the dampening effect of ESG rating disagreement on TFP. Dose this contradicts exiting researches that show firms tend to increase their green innovation efforts when ESG rating disagreement is present(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024b). This observation is very likely caused by a remedial action taken by firms in response to ESG rating disagreement, as Hou and Xie (2024) mentioned, such ESG rating disagreement is negatively correlated with the quality of corporate green innovation. According to stakeholder theory, firm’s remedial green innovations may be seen as "greenwashing" when there is disagreement over ESG ratings, which will lead to higher costs of debt financing for corporations, ultimately bringing about a decline in TFP(Peng and Xie, 2024).
Table 4: Green Innovation Mechanism
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.0264 -0.0205 -0.0199 -0.0167 -0.0130 -0.0136
(-1.2045) (-1.0114) (-0.9590) (-0.8367) (-0.7014) (-0.7069)
$EI$ 0.0412*** 0.0283*** 0.0253*** 0.0223*** 0.0142*** 0.0126***
(8.4064) (6.2383) (5.4522) (4.9596) (3.3848) (2.8967)
$Dis*EI$0.0366*** -0.0292*** -0.0296*** -0.0249** -0.0199** -0.0207**
(-3.1706) (-2.7348) (-2.7184) (-2.3693) (-2.0401) (-2.0432)
$TobinQ$0.0030 -0.0016 -0.0007
(-0.7895) (-0.4611) (-0.1778)
$Top10$$0$$0$0005
(-1.7413) (-1.7832) (-0.8682)
$ListAge$ 0.1012*** 0.0869*** 0.0172
(3.4856) (3.2139) (0.6122)
$Lev$0.1692*** -0.1550*** -0.1472***
(-3.2834) (-3.2318) (-2.9594)
$Size$ 0.3664*** 0.2537*** 0.2208***
(28.243) (21.017) (17.632)
$Avg$ 0.2297*** 0.2698*** 0.2673***
(22.666) (28.608) (27.331)
$Der$$0**$0190*** -0.0211***
(-3.1573) (-3.0646) (-3.2805)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
$N$ 13417 13417 13417 13417 13417 13417
$R^{2}$ 0.0092 0.0052 0.0042 0.1806 0.1680 0.1417
Figure 1: Marginal Effect
4.3.2 The Role of Financing Constraints
Undoubtedly, ESG investment plays a crucial role in promoting ESG development, making it essential to examine the impact of ESG rating disagreement on corporate financing. According to the information asymmetry theory discussed in the previous section, ESG rating disagreement introduces noise, potentially contributes financing constraints for firms. This study focuses exclusively on data from domestic ESG rating agencies in China, in contrast to Li et al. (2024), which also includes data from foreign rating agencies to explore this mechanism. Hence, we employ the WW and KZ indices form CSAMR to enhance the robustness of the empirical evidence. Furthermore, we follow Jiang (2022)’s research on the transmission mechanism, to avoid financing constraints being both exogenous and endogenous. We only use independent variable to regress with Finacing constraints, since numerous researches have demonstrated that financing constraints are detrimental to TFP improvement(Hopenhayn, 2014; Zhang et al., 2021; Piao et al., 2023).
Table 5 reports the regression results. The $Disb$ is the variable by one period. So we concludes that increased ESG rating disagreement reduces corporate financing constraints, thereby lowering TFP.
Table 5: Financing Constraints Mechanism
Variables WW
$WW$ $WW$ $KZ$ $KZ$ $WW$ $WW$ $KZ$ $KZ$
$Dis$ 0.0122*** 0.2166*** 0.0089*** 0.1815***
(4.7408) (3.0974) (3.6730) (2.6598)
$Disb$ 0.0122*** 0.2142*** 0.0074*** 0.1392**
(4.7691) (3.0604) (3.3430) (2.2547)
$TobinQ$0.0022*** -0.0006 0.2724*** 0.0237
(-3.9269) -0.9530 (19.528) (1.4433)
$Top10$0.0007*** -0.0006*** 0.0011 -0.0061**
(-8.8614) -6.9840 (0.4783) (-2.3496)
$ListAge$ 0.0204*** 0.0143*** 0.5530*** 1.2048***
(6.1906) (2.9544) (6.5333) (9.5882)
$Lev$ 0.0370*** 0.0359*** 6.7446*** 3.1453***
(5.0820) (4.5416) (32.842) (14.109)
$Size$0.0614*** -0.0455*** -0.4512*** -0.2274***
(-36.832) (-22.973) (-9.5683) (-4.0781)
$Avg$$0**$0113*** -0.5288*** -0.0107
(-18.835) (-7.3282) (-13.531) (-0.2432)
$Der$ 0.0030*** 0.0011 0.0191 -0.0027
(3.4646) (1.2994) (0.6820) (-0.0931)
Controls NO NO NO YES YES YES YES YES
Year YES YES YES YES YES YES YES YES
Id YES YES YES YES YES YES YES YES
$N$ 10059 10059 13417 13417 10059 10059 13417 13417
$R^{2}$ 0.0034 0.0034 0.0010 0.0010 0.2647 0.1221 0.2250 0.0517
  • Note: Similar to the robustness tests, $Dis$ corresponds to regressions on ESG rating disagreement in year t and financing constraints in year t+1, while $Disb$ corresponds to regressions on ESG rating disagreement in year t-1 and financing constraints in year t.
4.4 Heterogeneity Analysis
To further explore whether the dampening effect of ESG rating disagreement on TFP exhibits heterogeneity, this study conducts a subgroup regression analysis. Given that the sample size decreases after subgrouping, making the results more susceptible to the influence of outliers, the data was Winsorized at the 2% level on both ends before performing the subgroup regression and conducting a Chow Test on $Dis$.
Table 6 reports the results of the heterogeneity analysis. The results indicate that the effect is more significant in state-owned firms, non-capital-intensive firms, and low-pollution firms. The above heterogeneity may exist because, first, there is a greater willingness to engage in ESG construction among non-state-owned firms due to the fact that they tend to focus more on ESG performance to attract more investment; second, usually capital-intensive firms are less affected by market forces due to their own low variable costs and high fixed costs; third, in the case of highly polluting firms that are susceptible to a variety of environmental policies and regulations in their financing, such firms are less likely to invest in ESG performance, and therefore the impact of divergent ESG ratings is less significant.
Table 6: Heterogeneity Analysis
Variables TFP
State-Owned Non-State-Owned Capital-Intensive Non-Capital-Intensive High-Pollution Low-Pollution
$TFP_{LP}$
$Dis$0.0622** -$0$0322 -0.0462*** -0.0239 -0.0481***
(-2.0516) (-1.8294) (-1.1009) (-2.8884) (-1.0401) (-2.7959)
Chow Test 500.19*** 289.12*** 544.57***
$R^{2}$ 0.1573 0.1886 0.1741 0.1817 0.2010 0.1754
$TFP_{OP}$
$Dis$0.0553* -0.0268 -0.0097 -0.0458*** -0.0070 -0.0493**
(-1.8608) (-1.5208) (-0.3229) (-2.6338) (-0.3149) (-2.4735)
Chow Test 529.58*** 269.31*** 534.66***
$R^{2}$ 0.1759 0.1593 0.1672 0.1663 0.1930 0.1595
$TFP_{GMM}$
$Dis$0.0618** -$0$0278 -0.0471*** -0.0134 -0.0519**
(-2.1381) (-1.6745) (-0.9906) (-2.6889) (-0.6185) (-2.5484)
Chow Test 498.97*** 248.15*** 475.42***
$R^{2}$ 0.1607 0.1254 0.1504 0.1377 0.1691 0.1321
Controls YES YES YES YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
Double Clustering YES YES YES YES YES YES
$N$ 4416 9001 2342 11075 3900 9517
5 Further Discussion
In the above analysis we can get that ESG rating disagreement brings about a decrease in TFP, which is obtained based on linear regression. In order to further explore the relationship between ESG rating divergence and TFP under the nonlinear assumption, we consider a regression analysis using a machine learning approach that can be interpreted with SHapley Additive exPlanations(SHAP) introduced by Scott et al. (2017), which can observe whether ESG rating disagreement is crucial for explaining TFP under the assumption of nonlinear system and explore the nonlinear influence relationship between the two variables.
Following Xue et al. (2024), we use the XGBoost machine learning model introduced by Chen and Guestrin (2016) to predict $TFP_{LP}$. XGBoost is an ensemble learner based on gradient boosted trees (GBDT) that has demonstrated excellent performance in many applications, even outperforming deep neural networks in some tasks. In this paper, the Optuna library developed by Akiba et al. (2019), combed with early stopping and cross-validation strategies is employed for parameter selection. After that, we adopt the SHAP (SHapley Additive exPlanations) method proposed by Scott et al. (2017) to explain XGBoost. The variables include the ESG rating disagreement ($Dis$ $Disb$), all control variables, and institutional ownership percentage ($Inst$).
We use 80% of the dataset for training and 20% for testing and conducted three rounds of training: the first round included ESG rating disagreement, the second round added ESG rating disagreement from the previous period, and the third round did not include ESG rating disagreement. Table 7 reports the evaluation metrics for the three rounds of training. It can be observed that using $R^{2}$ as the evaluation metric, the model’s predictive performance improved by 4.3% and 6.23% respectively compared to the model without ESG rating disagreement, indicating that ESG rating disagreement aids in the short-term prediction of $TFP_{LP}$.
Table 7: Performance Metrics for Training and Test Sets
Metric Training Set 1 Training Set 2 Training Set 3 Test Set 1 Test Set 2 Test Set 3
MSE 0.2609 0.2274 0.2818 0.3492 0.3300 0.3835
RMSE 0.5108 0.4769 0.5309 0.5909 0.5744 00.6192
MAE 0.3995 0.3729 0.4145 0.4589 0.4523 0.4818
MAPE 4.7786% 4.4569% 4.9457% 5.4649% 5.3644% 5.7293%
R2 0.7685 0.7969 0.7500 0.7045 0.7176 0.6755
To analyze the influence of ESG rating disagreement on $TFP_{LP}$ in the predictive task, SHAP was used to interpret the XGBoost model. SHAP values, rooted in game theory, quantify each feature’s contribution to the model output by treating features as "players" in a game. The Shapley value $\phi_{i}$ for a feature $i$ is calculated as:
$\phi_{i}=\sum_{S\subseteq N\setminus\{i\}}\frac{|S|!\cdot(|N|-|S|-1)!}{|N|!}\left[f(S\cup\{i\})-f(S)\right]$ (3)
where $S$ is a subset excluding $i$, $f(S)$ is the model output for $S$, and $f(S\cup\{i\})$ is the output after adding $i$ to $S$.
In this study, based on the previous training results, we chose the model with one-period lagged ESG rating disagreement for our SHAP analysis, and the SHAP values were calculated for the test set. Figure 2 reports the feature importance, showing that the contribution ratio of $Disb$ is 0.14, again indicating that $Disb$ is crucial for predicting total factor productivity. Figure 3 reports the SHAP beeswarm plot, it can be observed that high ESG rating divergence suppresses a company’s total factor productivity, while in companies with low ESG ratings and small divergences, ESG rating disagreement is even positively correlated with TFP in the prediction task. The reason for this phenomenon is that the investment market is more tolerant of small ESG rating divergence, and listed companies will not face significant financing pressure due to small divergence.
Figure 2: Feature Importance
Figure 3: SHAP Beeswarm Plot
6 Conclusion
ESG rating disagreement poses challenges to the enhancement of corporate TFP. This study utilizes financial data from China’s A-share listed companies and ESG rating information from domestic ESG rating agencies between 2015 and 2022 to explore the relationship between ESG rating disagreements and firms’ TFP. The findings robustly indicate that ESG rating disagreements have a negative impact on firms’ TFP. The mechanism analysis reveals that green innovation instead strengthens the dampening effect of ESG rating disagreement on TFP; moreover, ESG rating disagreement strengthens firms’ financing constraints, thereby hindering the improvement of TFP. Heterogeneity analysis further indicates that the negative impact of ESG rating disagreements on TFP is more pronounced in state-owned enterprises, non-capital-intensive firms, and companies with lower levels of pollution. Additionally, an important finding from the analysis using machine learning methods suggests that ESG rating disagreements serve as a significant indicator for short-term predicting firms’ TFP, providing new insights for companies to engage in ESG risk management to enhance TFP.
Appendix A Appendix
A.1 Robustness Tests
Table 8: Robustness Tests (I)
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.0662*** -0.0524** -0.0524** -0.0246** -0.0159** -0.0180**
(-2.7472) (-2.2660) (-2.2781) (-2.2085) (-2.4677) (-2.0769)
$TobinQ$ 0.0218*** 0.0141*** 0.0150***
(6.8126) (7.0455) (6.1013)
$Top10$ 0.0005 0.0002 0.0004
(0.8363) (0.5950) (0.8865)
$ListAge$$0*$0291** -0.0678***
(-2.3362) (-2.2563) (-4.0705)
$Lev$ 0.1415 0.0739 0.0650
(1.8712) (1.5504) (1.2862)
$Size$ 0.4212*** 0.2448*** 0.1800***
(21.576) (18.539) (11.207)
$Avg$ 0.7135*** 0.8172*** 0.8193***
(38.466) (64.699) (55.668)
$Der$0.0257** -0.0113 -0.0135
(-2.0169) (-1.3877) (-1.5955)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
Doubel Clustering YES YES YES YES YES YES
$N$ 13417 13417 13417 13417 13417 13417
$R^{2}$ 0.0015 0.0011 0.0011 0.7027 0.8496 0.7734
Table 9: Robustness Tests (II)
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.0647** -0.0519** -$0*$0419*** -0.0342** -0.0346**
(-2.5465) (-1.9984) (-2.0263) (-2.6444) (-1.9838) (-1.9658)
$TobinQ$0.0027 -0.0013 -0.0003
(-0.4030) (-0.2619) (-0.0725)
$Top10$0.0009 -0.0009 -0.0005
(-0.9777) (-0.9697) (-0.5039)
$ListAge$ 0.0906** 0.0796** 0.0058
(2.4895) (2.3720) (0.1639)
$Lev$0.1555 -0.1408 -0.1334
(-1.3079) (-1.2639) (-1.1864)
$Size$ 0.3762*** 0.2596*** 0.2262***
(14.285) (13.442) (10.036)
$Avg$ 0.2277*** 0.2680** 0.2660**
(2.6951) (2.7597) (2.6807)
$Der$$0$0194 -0.0217
(-1.9537) (-1.7844) (-1.9273)
$N.TFP_{province}$ 0.0993 0.0847 0.0403
(0.6302) (0.5857) (0.2531)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
Doubel Clustering YES YES YES YES YES YES
$N$ 12893 12893 12893 12893 12893 12893
$R^{2}$ 0.0014 0.0011 0.0010 0.1793 0.1670 0.1406
Table 10: Robustness Tests (III)
Variables Dependent Variable
$Dis$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$IV(\hat{Dis})$0.3336*** -0.0479*** -0.0408** -0.0424**
(-602.14) (-3.0748) (-2.3966) (-2.3533)
$TobinQ$0.0007 -0.0027 -0.0013 -0.0004
(-0.5971) (-0.4037) (-0.2634) (-0.0756)
$Top10$ 3.476e-05 -0.0010 -0.0009 -0.0005
(0.3084) (-0.9808) (-0.9708) (-0.5052)
$ListAge$0.0056 0.0910** 0.0799** 0.0061
(-0.9005) (2.4981) (2.3788) (0.1734)
$Lev$0.0029 -0.1554 -0.1407 -0.1333
(-0.3495) (-1.3058) (-1.2618) (-1.1844)
$Size$0.0045** 0.3760*** 0.2594*** 0.2260***
(-2.6017) (14.301) (13.453) (10.029)
$Avg$0.0017 0.2278*** 0.2680** 0.2660**
(-0.7314) (2.6949) (2.7600) (2.6811)
$Der$ 0.0037** -$0$0193 -0.0216
(2.8094) (-1.9412) (-1.7701) (-1.9119)
$N.TFP_{province}$ 0.0100 0.0999 0.0851 0.0408
(0.5583) (0.6350) (0.5902) (0.2565)
Controls YES YES YES YES
Year YES YES YES YES
ID YES YES YES YES
Double Clustering YES YES YES YES
$F$ 2.983e+04 227.22 208.52 170.30
$N$ 16907 12892 12892 12892
R2 0.9542 0.1794 0.1671 0.1408
  • Note: The second stage of 2SLS regression follows the baseline regression with a one-period lag.
Table 11: Robustness Tests (IV)
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.1054*** -0.0959*** -0.0908*** -$0*$0676** -0.0645**
(-3.0359) (-2.9849) (-2.7841) (-2.3040) (-2.2900) (-2.1303)
$TobinQ$ 0.0007 0.0002 -0.0026
(0.0934) (0.0351) (-0.3889)
$Top10$ 0.0005 0.0004 0.0007
(0.5206) (0.3886) (0.6853)
$ListAge$ 0.1119 0.1001 0.0149
(1.6127) (1.5470) (0.2245)
$Lev$0.0295 -0.0270 -0.0174
(-0.3276) (-0.3216) (-0.2021)
$Size$ 0.3899*** 0.2685*** 0.2291***
(16.968) (12.538) (10.440)
$Avg$ 0.2343*** 0.2761*** 0.2832***
(13.111) (16.582) (16.598)
$Der$0.0408*** -0.0368*** -0.0385***
(-3.4013) (-3.2928) (-3.3577)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
ID YES YES YES YES YES YES
Double Clustering NO NO NO NO NO NO
$N$ 5368 5368 5368 5368 5368 5368
R2 0.0028 0.0027 0.0024 0.1739 0.1628 0.1455
A.2 Optuna Parameter Range & Correlation Coefficient
Table 12: Correlation Matrix
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $Dis$ $TobinQ$ $Top10$ $ListAge$ $Lev$ $Size$ $Avg$ $Der$ $N.TFP_{province}$ $IV$
$TFP_{LP}$ 1.000 0.951 0.918 -0.066 -0.200 0.127 0.329 0.444 0.801 0.648 0.344 -0.037 0.082
$TFP_{OP}$ 0.951 1.000 0.976 -0.054 -0.183 0.104 0.303 0.405 0.707 0.841 0.325 -0.036 0.067
$TFP_{GMM}$ 0.918 0.976 1.000 -0.041 -0.132 0.080 0.244 0.356 0.583 0.823 0.295 -0.005 0.059
$Dis$0.066 -0.054 -0.041 1.000 0.014 -0.078 0.025 0.039 -0.071 -0.024 0.052 -0.015 -0.974
$TobinQ$0.200 -0.183 -0.132 0.014 1.000 -0.036 -0.094 -0.249 -0.283 -0.116 -0.173 0.012 -0.008
$Top10$ 0.127 0.104 0.080 -0.078 -0.036 1.000 -0.364 -0.066 0.147 0.044 -0.052 0.038 0.087
$ListAge$ 0.329 0.303 0.244 0.025 -0.094 -0.364 1.000 0.292 0.424 0.204 0.240 -0.184 -0.024
$Lev$ 0.444 0.405 0.356 0.039 -0.249 -0.066 0.292 1.000 0.458 0.256 0.843 -0.044 -0.038
$Size$ 0.801 0.707 0.583 -0.071 -0.283 0.147 0.424 0.458 1.000 0.413 0.359 -0.104 0.082
$Avg$ 0.648 0.841 0.823 -0.024 -0.116 0.044 0.204 0.256 0.413 1.000 0.222 -0.041 0.027
$Der$ 0.344 0.325 0.295 0.052 -0.173 -0.052 0.240 0.843 0.359 0.222 1.000 -0.047 -0.046
$N.TFP_{province}$0.037 -0.036 -0.005 -0.015 0.012 0.038 -0.184 -0.044 -0.104 -0.041 -0.047 1.000 0.014
$IV$ 0.082 0.067 0.059 -0.974 -0.008 0.087 -0.024 -0.038 0.082 0.027 -0.046 0.014 1.000
Table 13: Hyperparameters for XGBoost Model
Parameter Range(optuna) / Value
loss_function rmse
learning_rate [0.005, 0.2] (log scale)
max_depth [1, 3]
min_child_weight [1, 5] (log scale)
subsample [0.5, 0.8]
colsample_bytree [0.5, 0.8]
reg_lambda [1, 10] (log scale)
reg_alpha [1, 10] (log scale)
tree_method hist
device cuda
num_boost_round 1000
nfold 20
early_stopping_rounds 5
seed 42