### v5

# ESG Rating Disagreement and Corporate Total Factor Productivity: Inference and Prediction

Abstract
This paper examines how ESG rating disagreement ($Dis$) affects corporate total factor productivity (TFP) in China based on data of A-share listed companies from 2015 to 2022. We find that $Dis$ reduces TFP, especially in state-owned, non-capital-intensive, low-pollution and high-tech firms, green innovation strengthens the dampening effect of $Dis$ on TFP, and that $Dis$ lowers corporate TFP by increasing financing constraints and weakening human capital. Furthermore, XGBoost regression demonstrates that $Dis$ plays a significant role in predicting TFP, with SHAP showing that the dampening effect of ESG rating disagreement on TFP is still pronounced in firms with large $Dis$ values.
JEL Codes: G32, G24, C54.
Keywords: ESG rating disagreement; Total factor productivity; Green innovation; Financing constraints; Human capital; Machine learning
1 Introduction
Since 1987, the United Nations has championed sustainable development and introduced ESG (Environmental, Social, and Governance) in 2004. Over time, these principles have become key components of global investment strategies and corporate assessments. In recent years, China has embraced sustainable development through initiatives like the philosophy “green mountains and clear waters are as valuable as mountains of gold and silver.” On December 27, 2023, the Central Committee of the Communist Party of China and the State Council issued the Guidelines for the Comprehensive Advancement of Beautiful China Construction, which emphasized creating an ESG Evaluation System with Chinese Characteristics.
Despite recent ESG policy efforts, significant disagreements persist within ESG rating systems. Since ESG ratings were introduced in China, multiple agencies have issued their own reports, often resulting in substantial discrepancies. Such differences have considerable implications for evaluated companies. Recent studies highlight the economic impacts of ESG rating disagreement: stock returns suffer (Wang et al., 2024; Zhao et al., 2024); analyst forecast errors rise (Liu et al., 2024); green innovation encouragement (Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); real earnings management increases in both the short and long term (Li and Cheng, 2024); firms with consistently high ratings benefit from lower costs and increased debt financing (Guo et al., 2024); disagreement poses a risk of stock price collapse (Luo et al., 2023); audit fees increase due to information asymmetry (Ling et al., 2024); and market pressures on managers intensify (Chen and Wang, 2025).
However, most recent studies have not sufficiently examined ESG rating disagreement’s impact on a company’s overall development. To address this gap, we focus on TFP as an indicator of high-quality development in this study.
2 Theoretical Analysis and Research Hypotheses
ESG rating disagreement can confuse stakeholders and increase information asymmetry between companies and investors. While ESG ratings help reduce information asymmetry by providing non-financial insights(Xiaohong and Zhenghan, 2023), disagreement can amplify asymmetry, creating noise that misleads investors(Berg et al., 2022), raises financing constraints, and ultimately lowers TFP. Stakeholder theory suggests that companies may respond to disagreement by pursuing green innovation(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a), but this can appear as “greenwashing” to some stakeholders, undermining financing and reducing TFP(Peng and Xie, 2024). Disagreement may also prompt firms to prioritize ESG recognition, causing tension between external investors and internal employees, potentially leading to a human capital outflow and further TFP decline.
Hence, we propose the research hypothesis H1: ESG rating disagreement reduces corporate TFP.
3 Data and Empirical Design
3.1 Sample Selection
We use financial data of A-share listed companies in China from 2015 to 2022111Prior 2015, the widely recognized local ESG rating agency in China was Huazheng., along with ESG ratings from Huazheng, Wind, SynTao Green Finance, and Susallwave. The data processing steps are as follows: (1) exclude companies in the financial and real estate sectors; (2) exclude ST, *ST, and delisted firms; (3) exclude companies with only one ESG rating in a year; (4) exclude firms with missing key indicators; (5) apply Winsorization at the 1% level on both tails. The data comes from the CSMAR and CNRDS databases and the four rating agencies, yielding 3656 firms and 13,417 samples in the baseline regression.
3.2 Model Specification
To test the hypotheses presented earlier,we conduct the following baseline regression model [^2] Considering the time required for information to be transmitted and for the market to react and mitigate potential reverse causation problems, TFP is lagged by one period.:
$TFP_{i,t+1}=\alpha_{0}+\alpha_{1}Dis_{i,t}+\alpha_{2}Control_{i,t}+Year_{t}+Id_{i}+\epsilon_{i,t}$ (1)
where $TFP_{i,t+1}$ is the total factor productivity of company $i$ in year $t+1$, $Dis_{i,t}$ is the ESG rating discrepancy, $Control_{i,t}$ represents the control variables, $Year_{t}$ and $Id_{i}$ denote year and company fixed effects, respectively, and $\epsilon_{i,t}$ is the error term.
3.3 Variable Descriptions
(1) Dependent Variable
The dependent variable in this study is Corporate Total Factor Productivity. Considering the robustness of the empirical process, we use three methods to measure TFP, they are LP introduced byLevinsohn and Petrin (2003) and adopted by Lu and Lian (2012), OP by Olley and Pakes (1992) and GMM by Blundell and Bond (1998). In the prediction task, we use TFP measured by LP method as the dependent variable(Xue et al., 2024).
(2) Independent Variable
The independent variable in this study is ESG rating disagreement. Domestic ESG rating agencies have advantages in terms of time and spatial proximity that foreign agencies cannot match, making them more likely to influence the behavior of domestic firms. Therefore, this study uses ESG rating data from four Chinese agencies: Huazheng, Wind, SynTao Green Finance, and Susallwave. Following Avramov et al. (2022), the ratings were assigned positive values according to their rating levels, with a difference of 1 between levels. We define the rankings as the values normalized to the range [0, 1] for the given rating agency in a given year, and the pairwise standard deviation of these rankings is used as the measure of $Dis$.
$Dis_{i,t}=\sqrt{\frac{1}{n-1}\sum_{k\neq j}^{n}(rank_{i,t,k}-rank_{i,t,j})^{2}}$ (2)
where $rank_{i,j,k}$ represents the normalized ranking of corporate $i$ in year $t$ given by agency $k$, $n$ represents the number of pairs, $k$ and $j$ represent different ESG rating agencies.
(3) Control Variables
Based on the characteristics of our research and with reference to the works of Yang et al. (2024) and Xue et al. (2024), the control variables include firm value, top 10 shareholders’ ownership percentage, years listed, leverage ratio, firm size, revenue per capita, and ownership ratio. The main variables are defined in Table 1.
[Insert Table 1 Here]
4 Empirical Analysis
4.1 Baseline Regression
Table 2 reports the results of the baseline regression, with clustering at the individual and time levels. The results indicate that ESG rating disagreement reduces corporate TFP, providing preliminary support for the validity of Hypothesis H1.
[Insert Table 2 Here]
4.2 Mechanism Analysis
4.2.1 The Role of Green Innovation
Myriad existing researches demonstrate that ESG rating disagreement can drive firms to enhance green innovation(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024b). However, does this green innovation truly contribute to substantial corporate development? This study introduces green innovation as a mechanism variable, measured by the natural logarithm of the number of green patents held by a company. To examine the role of green innovation in the relationship between ESG rating disagreement and TFP, an interaction term between green innovation and ESG rating disagreement is constructed, with other settings remaining consistent with Equation 1.
Table 3 and Figure 1 report the results. The results show that when there are no ESG rating discrepancies, green innovation promotes the improvement of TFP. Conversely, when there are ESG rating discrepancies, it strengthens the dampening effect of ESG rating disagreement on TFP. The result seem contradict exiting researches, which argue ESG rating disagreement can drive firms to enhance green innovation(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024b). This observation is very likely caused by a remedial action taken by firms in response to ESG rating disagreement, as Hou and Xie (2024) mentioned, such ESG rating disagreement is negatively correlated with the quality of corporate green innovation. According to stakeholder theory, firm’s remedial green innovations may be seen as "greenwashing" when there is disagreement over ESG ratings, which will lead to higher costs of debt financing for corporations, ultimately bringing about a decline in TFP(Peng and Xie, 2024).
[Insert Table 3 Here]
[Insert Figure 1 Here]
To further investigate whether the moderating effect varies across levels of ESG rating disagreement, we conducted regressions separately for the upper and lower 50% quartiles of ESG rating disagreement, and the Table 4 show that the moderating effect is significant only when the ESG rating disagreement is high, further supporting the notion that substantial ESG rating disagreement may lead to stakeholder distrust in firms’ green innovation.
[Insert Table 4 Here]
4.2.2 The Role of Financing Constraint
Undoubtedly, ESG investment plays a crucial role in promoting ESG development, making it essential to examine the impact of ESG rating disagreement on corporate financing. According to the information asymmetry theory discussed in the previous section, ESG rating disagreement introduces noise, potentially contributes financing constraints for firms. This study focuses exclusively on data from domestic ESG rating agencies in China, in contrast to Li et al. (2024), which also includes data from foreign rating agencies to explore this mechanism. Hence, we employ the WW and KZ indices form CSAMR to enhance the robustness of the empirical evidence. Furthermore, we follow Jiang (2022)’s research on the transmission mechanism, to avoid financing constraints being both exogenous and endogenous. We only use independent variable to regress with finacing constraints, since numerous researches have demonstrated that financing constraints are detrimental to TFP improvement(Hopenhayn, 2014; Zhang et al., 2021; Piao et al., 2023).
Table 5 reports the regression results. The $Disb$ is the variable by one period. So we concludes that increased ESG rating disagreement strengthens corporate financing constraints, thereby lowering TFP.
[Insert Table 5 Here]
4.2.3 The Role of Human Capital
Human capital is widely acknowledged as a positive contributor to TFP(Miller and Upadhyay, 2000; Xu et al., 2008; Huang et al., 2024). Employees with a bachelor’s or master’s degree typically reflects higher levels of knowledge, skills, and expertise among workers, enabling them to perform task more effciently, reduce errors, and enhance productivity. In this subsection, we delve deeper into how ESG rating disagreement interacts with human capital and its potential impact on a company’s TFP, with a specific focus on employees holding bachelor’s or master’s degrees. Building on Jiang (2022)’s research on the transmission mechanism, and following a similar approach to verifying the financing constraint mechanism to avoid endogeneity issues in mediating variables, we perform regressions of $Dis$ on $Master$ and $Dis$ on $Undergrad$.
Table 6 reports the regression results, showing that Dis is detrimental to the development of firms’ human capital, which serves as a crucial driver of TFP. Based on this evidence, we argue that Dis reduces TFP by impairing human capital.
[Insert Table 6 Here]
4.3 Heterogeneity Analysis
To further explore whether the dampening effect of ESG rating disagreement on TFP exhibits heterogeneity, this study conducts a subgroup regression analysis. Given that the sample size decreases after subgrouping, making the results more susceptible to the influence of outliers, the data is Winsorized at the 2% level on both ends before performing the subgroup regression and conducting a Chow Test on $Dis$.
Table 7 reports the results of the heterogeneity analysis. The results indicate that the effect is more significant in state-owned firms, non-capital-intensive firms, low-pollution firms and high-tech firms. This heterogeneity may arise for the following reasons: first, non-state-owned firms are generally more willing to engage in ESG initiatives as they tend to focus more on ESG performance to attract investment; second, capital-intensive firms are usually less influenced by market forces due to their low variable costs and high fixed costs, which buffer them from the pressures of ESG-related changes; third, highly polluting firms, which are often subject to stringent environmental policies and regulations in their financing, are less likely to invest in ESG performance, making the impact of divergent ESG ratings less pronounced. And, finally, as previously analyzed, ESG rating disagreements can lead to the loss of highly qualified employees, on whom high-tech firms are particularly dependent. Consequently, the negative impact of disagreements on TFP is more significant for high-tech firms.
[Insert Table 7 Here]
5 Further Discussion
In the above analysis we can see that ESG rating disagreement brings about a decrease in TFP, which is obtained based on linear regression. In order to further explore the relationship between ESG rating divergence and TFP under the nonlinear assumption, we consider a regression analysis using a machine learning approach that can be interpreted with SHapley Additive exPlanations(SHAP) introduced by Scott et al. (2017), which can observe whether ESG rating disagreement is crucial for explaining TFP under the assumption of nonlinear system and explore the nonlinear influence relationship between the two variables.
Following Xue et al. (2024), we use the XGBoost machine learning model introduced by Chen and Guestrin (2016) to predict $TFP_{LP}$. XGBoost is an ensemble learner based on gradient boosted trees (GBDT) that has demonstrated excellent performance in many applications, even outperforming deep neural networks in tabular data(Grinsztajn et al., 2022). In this paper, the Optuna library developed by Akiba et al. (2019), combed with early stopping and cross-validation strategies is employed for parameter selection. After that, we adopt the SHAP (SHapley Additive exPlanations) method proposed by Scott et al. (2017) to explain XGBoost. The variables include the ESG rating disagreement ($Dis$ $Disb$), all control variables, and institutional ownership percentage ($Inst$).
We use 80% of the dataset for training and 20% for testing, conducting three rounds of training. The first round includes ESG rating disagreement (Dis), the second round incorporates ESG rating disagreement from the previous period, and the third round excludes ESG rating disagreement altogether. Table 8 presents the evaluation metrics for these three rounds of training.
First, compared to the linear model (Set4), XGBoost demonstrates significantly better performance across all metrics, highlighting the machine learning model’s ability to capture nonlinear relationships. Second, when comparing XGBoost without incorporating Dis (Set3), the model’s performance improves by 4.3% (Set1) and 6.23% (Set2), respectively, using $R^{2}$ as the evaluation criterion. This improvement underscores the predictive value of ESG ratings in modeling TFP.
[Insert Table 8 Here]
To analyze the influence of ESG rating disagreement on $TFP_{LP}$ in our predictive task, we use SHAP to interpret the results of the XGBoost model. SHAP values are grounded in cooperative game theory and provide a way to understand each feature’s contribution to the model output by treating features as "players" in a game. The contribution, or Shapley value $\phi_{i}$, for a particular feature $i$ is computed by averaging its marginal contributions across all possible combinations of features.
Mathematically, the Shapley value $\phi_{i}$ for feature $i$ is defined as:
$\phi_{i}=\sum_{S\subseteq N\setminus\{i\}}\frac{|S|!\cdot(|N|-|S|-1)!}{|N|!}\left[f(S\cup\{i\})-f(S)\right]$ (3)
where $N$ represents the set of all features, $S\subseteq N\setminus\{i\}$ is a subset excluding feature $i$, $|S|$ denotes the cardinality of subset $S$, $f(S)$ indicates the model output using feature subset $S$, and $f(S\cup\{i\})$ represents the output when feature $i$ is added to $S$.
The term $\frac{|S|!\cdot(|N|-|S|-1)!}{|N|!}$ in the equation represents a weighting factor that accounts for all possible orders in which features can be added to the model. By calculating $\phi_{i}$ across all possible subsets $S$, SHAP captures the average marginal effect of adding feature $i$ to different subsets of features, thereby quantifying the feature’s contribution to the model’s output.
In this study, based on the previous training results, we chose the model with one-period lagged ESG rating disagreement for our SHAP analysis, and the SHAP values were calculated for the test set. Figure 2 reports the feature importance, showing that the contribution ratio of $Disb$ is 0.14, again indicating that $Disb$ is crucial for predicting TFP. Figure 3 reports the SHAP beeswarm plot, it can be observed that high ESG rating divergence suppresses a company’s TFP, while in companies with low ESG ratings and small divergences, ESG rating disagreement is even positively correlated with TFP in the prediction task. The reason for this phenomenon is that the investment market is more tolerant of small ESG rating divergence, and listed companies will not face significant financing pressure due to small divergence.
[Insert Figure 2 Here]
[Insert Figure 3 Here]
6 Conclusion and Implication
ESG rating disagreement presents significant challenges to the improvement of corporate TFP, underscoring the urgent need for firms to address this issue. Mechanism analysis reveals that green innovation can exacerbate the negative effects of ESG rating disagreement on TFP. Additionally, these disagreement hinder TFP growth by intensifying financing constraints and weakening human capital. To mitigate these impacts, companies should focus on improving the quality of ESG information disclosure, carefully navigate the trade-offs between green innovation, financing needs, and strive to balance the interests of diverse stakeholders. Regulatory authorities, in turn, should strengthen the scrutiny of green innovation patents to ensure their quality and prevent superficial compliance efforts.
Heterogeneity analysis demonstrates that the negative effects of ESG rating disagreement on TFP are more pronounced in state-owned enterprises, non-capital-intensive firms, low-pollution industries, and high-tech enterprises. These entities should pay closer attention to ESG rating discrepancies, while regulators should provide targeted support to low-pollution and high-tech firms to foster sustainable development and technological innovation. Moreover, ESG rating disagreement serve as a critical predictor of TFP, with SHAP values highlighting their substantial and ongoing negative impact. This indicator can be leveraged by listed companies and regulatory agencies as a monitoring and early warning tool to identify potential risks to TFP and implement proactive measures.
We believe it would be worthwhile to further investigate how green innovation amplifies the negative impact of ESG rating divergence on TFP, as well as to integrate SHAP with microeconometric methods for more robust empirical analysis.
Appendix A Online Appendix
A.1 Descriptive Statistics
Table 9: Descriptive Statistics of Main Variables
Variables Count Mean SD Min 25% 50% 75% Max
$TFP_{LP}$ 13417 8.51 1.04 6.38 7.79 8.40 9.14 11.28
$TFP_{OP}$ 13417 6.83 0.85 5.14 6.24 6.72 7.33 9.17
$TFP_{GMM}$ 13417 5.72 0.80 4.09 5.18 5.63 6.18 8.12
$Dis$ 13417 0.26 0.19 0.01 0.11 0.24 0.34 0.83
$TobinQ$ 13417 2.03 1.32 0.83 1.22 1.61 2.31 8.34
$Top10$ 13417 56.77 15.15 23.59 45.91 56.95 67.87 90.89
$ListAge$ 13417 2.34 0.68 1.10 1.79 2.40 3.00 3.37
$Lev$ 13417 0.43 0.19 0.06 0.28 0.43 0.57 0.88
$Size$ 13417 22.52 1.34 20.09 21.54 22.32 23.31 26.41
$Avg$ 13417 4.79 0.80 3.07 4.25 4.69 5.25 7.17
$Der$ 13417 1.05 1.07 0.07 0.39 0.74 1.31 6.75
$EI$ 13417 1.30 1.49 0.00 0.00 1.10 2.30 7.74
$WW$ 10059 -1.03 0.08 -1.24 -1.08 -1.03 -0.98 -0.86
$KZ$ 13417 1.17 1.94 -4.35 0.02 1.36 2.49 6.22
$Master$ 10461 4.31 1.55 1.10 3.18 4.20 5.31 8.42
$Undergrad$ 12595 6.31 1.35 3.56 5.35 6.17 7.15 10.05
  • Note: The data in the table are descriptive statistics for deleting missing values after lagging one period of TFP.
A.2 Correlation Coefficient
Table 10: Correlation Matrix
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $Dis$ $TobinQ$ $Top10$ $ListAge$ $Lev$ $Size$ $Avg$ $Der$ $N.TFP_{province}$ $IV$
$TFP_{LP}$ 1.000 0.951 0.918 -0.066 -0.200 0.127 0.329 0.444 0.801 0.648 0.344 -0.037 0.082
$TFP_{OP}$ 0.951 1.000 0.976 -0.054 -0.183 0.104 0.303 0.405 0.707 0.841 0.325 -0.036 0.067
$TFP_{GMM}$ 0.918 0.976 1.000 -0.041 -0.132 0.080 0.244 0.356 0.583 0.823 0.295 -0.005 0.059
$Dis$0.066 -0.054 -0.041 1.000 0.014 -0.078 0.025 0.039 -0.071 -0.024 0.052 -0.015 -0.974
$TobinQ$0.200 -0.183 -0.132 0.014 1.000 -0.036 -0.094 -0.249 -0.283 -0.116 -0.173 0.012 -0.008
$Top10$ 0.127 0.104 0.080 -0.078 -0.036 1.000 -0.364 -0.066 0.147 0.044 -0.052 0.038 0.087
$ListAge$ 0.329 0.303 0.244 0.025 -0.094 -0.364 1.000 0.292 0.424 0.204 0.240 -0.184 -0.024
$Lev$ 0.444 0.405 0.356 0.039 -0.249 -0.066 0.292 1.000 0.458 0.256 0.843 -0.044 -0.038
$Size$ 0.801 0.707 0.583 -0.071 -0.283 0.147 0.424 0.458 1.000 0.413 0.359 -0.104 0.082
$Avg$ 0.648 0.841 0.823 -0.024 -0.116 0.044 0.204 0.256 0.413 1.000 0.222 -0.041 0.027
$Der$ 0.344 0.325 0.295 0.052 -0.173 -0.052 0.240 0.843 0.359 0.222 1.000 -0.047 -0.046
$N.TFP_{province}$0.037 -0.036 -0.005 -0.015 0.012 0.038 -0.184 -0.044 -0.104 -0.041 -0.047 1.000 0.014
$IV$ 0.082 0.067 0.059 -0.974 -0.008 0.087 -0.024 -0.038 0.082 0.027 -0.046 0.014 1.000
  • Note: This is the matrix of linear correlation coefficients for all variables.
A.3 Robustness Tests and Endogeneity Treatments
To ensure the robustness of the baseline regression results, the following robustness tests were conducted.
First, to verify whether Hypothesis H1 holds when only leading $Dis$ by one period, we lead the it by one period(see Table 11).
Second, since the TFP of firms in this study is influenced by regional productivity levels, to mitigate endogeneity issues arising from omitted variable bias, we include the provincial-level New quality productivity measured by LU et al. (2024) as a control variable (see Table 12).
Third, to further mitigate everse causality endogeneity issues, following Yang et al. (2022), we calculate the industry average of ESG rating disagreements by year and industry, and use the cube of the difference between the industry average and the ESG rating disagreements as an instrumental variable. A two-stage least squares (2SLS) estimation is conducted, and the Kleibergen-Paap rk LM statistics verify the verify the validity of instrumental variables(see Table 13).
Fourth, to address sample selection bias, we denote disagreements under 20% as non-existent and the remainder as present. Propensity scores are computed via logistic regression, followed by 1:1 nearest-neighbor matching and subsequent regression (see Table 14).
Table 11: Robustness Tests (I)
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.0662*** -0.0524** -0.0524** -0.0246** -0.0159** -0.0180**
(-2.7472) (-2.2660) (-2.2781) (-2.2085) (-2.4677) (-2.0769)
$TobinQ$ 0.0218*** 0.0141*** 0.0150***
(6.8126) (7.0455) (6.1013)
$Top10$ 0.0005 0.0002 0.0004
(0.8363) (0.5950) (0.8865)
$ListAge$$0*$0291** -0.0678***
(-2.3362) (-2.2563) (-4.0705)
$Lev$ 0.1415 0.0739 0.0650
(1.8712) (1.5504) (1.2862)
$Size$ 0.4212*** 0.2448*** 0.1800***
(21.576) (18.539) (11.207)
$Avg$ 0.7135*** 0.8172*** 0.8193***
(38.466) (64.699) (55.668)
$Der$0.0257** -0.0113 -0.0135
(-2.0169) (-1.3877) (-1.5955)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
Doubel Clustering YES YES YES YES YES YES
$N$ 13417 13417 13417 13417 13417 13417
$R^{2}$ 0.0015 0.0011 0.0011 0.7027 0.8496 0.7734
  • Note: This is the result of only front-loading $Dis$ compared to $TFP$ by one period. In this regression result, $Dis$ denotes ESG Rating Disagreement of $t-1$ year and $TFP$ denotes total factor productivity of $t$ year. Unless otherwise noted, $Dis$ and $TFP$ are consistent with the baseline regression settings.
Table 12: Robustness Tests (II)
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.0647** -0.0519** -$0*$0419*** -0.0342** -0.0346**
(-2.5465) (-1.9984) (-2.0263) (-2.6444) (-1.9838) (-1.9658)
$TobinQ$0.0027 -0.0013 -0.0003
(-0.4030) (-0.2619) (-0.0725)
$Top10$0.0009 -0.0009 -0.0005
(-0.9777) (-0.9697) (-0.5039)
$ListAge$ 0.0906** 0.0796** 0.0058
(2.4895) (2.3720) (0.1639)
$Lev$0.1555 -0.1408 -0.1334
(-1.3079) (-1.2639) (-1.1864)
$Size$ 0.3762*** 0.2596*** 0.2262***
(14.285) (13.442) (10.036)
$Avg$ 0.2277*** 0.2680** 0.2660**
(2.6951) (2.7597) (2.6807)
$Der$$0$0194 -0.0217
(-1.9537) (-1.7844) (-1.9273)
$N.TFP_{province}$ 0.0993 0.0847 0.0403
(0.6302) (0.5857) (0.2531)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
Doubel Clustering YES YES YES YES YES YES
$N$ 12893 12893 12893 12893 12893 12893
$R^{2}$ 0.0014 0.0011 0.0010 0.1793 0.1670 0.1406
  • Note: This is the result of the regression with the addition of control variables, and the other settings are consistent with the baseline regression. $N.TFP_{province}$ represents the level of total factor productivity at the provincial level.
Table 13: Robustness Tests (III)
Variables Dependent Variable
$Dis$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$IV(\hat{Dis})$0.3302*** -0.0476*** -0.0409** -0.0425**
(-328.74) (-3.0292) (-2.3843) (-2.3456)
$TobinQ$ 0.0002 -0.0027 -0.0013 -0.0003
(0.1447) (-0.3986) (-0.2574) (-0.0685)
$Top10$ 0.0001 -0.0009 -0.0009 -0.0005
(1.1390) (-0.9769) (-0.9679) (-0.5016)
$ListAge$0.0118 0.0906** 0.0797** 0.0058
(-1.2204) (2.4886) (2.3696) (0.1655)
$Lev$0.0061 -0.1555 -0.1408 -0.1335
(-0.7749) (-1.3069) (-1.2627) (-1.1853)
$Size$0.0044** 0.3761*** 0.2594*** 0.2260***
(-2.0583) (14.306) (13.451) (10.030)
$Avg$0.0015 0.2277*** 0.2680** 0.2660**
(-0.5293) (2.6949) (2.7600) (2.6812)
$Der$ 0.0045** -$0$0193 -0.0216
(3.1346) (-1.9398) (-1.7686) (-1.9099)
$N.TFP_{province}$0.0030 0.0994 0.0847 0.0404
(-0.2042) (0.6320) (0.5877) (0.2540)
Controls YES YES YES YES
Year YES YES YES YES
ID YES YES YES YES
Double Clustering YES YES YES YES
$F$ 2.049e+04 227.20 208.52 170.30
$N$ 12892 12892 12892 12892
Kleibergen-Paap rk LM 622.33****
R2 0.9517 0.1794 0.1671 0.1408
  • Note: This is the result of an instrumental variable test where $TFP$ is the t+1 period, just like the baseline regression.
Table 14: Robustness Tests (IV)
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.1054*** -0.0959*** -0.0908*** -$0*$0676** -0.0645**
(-3.0359) (-2.9849) (-2.7841) (-2.3040) (-2.2900) (-2.1303)
$TobinQ$ 0.0007 0.0002 -0.0026
(0.0934) (0.0351) (-0.3889)
$Top10$ 0.0005 0.0004 0.0007
(0.5206) (0.3886) (0.6853)
$ListAge$ 0.1119 0.1001 0.0149
(1.6127) (1.5470) (0.2245)
$Lev$0.0295 -0.0270 -0.0174
(-0.3276) (-0.3216) (-0.2021)
$Size$ 0.3899*** 0.2685*** 0.2291***
(16.968) (12.538) (10.440)
$Avg$ 0.2343*** 0.2761*** 0.2832***
(13.111) (16.582) (16.598)
$Der$0.0408*** -0.0368*** -0.0385***
(-3.4013) (-3.2928) (-3.3577)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
ID YES YES YES YES YES YES
Double Clustering NO NO NO NO NO NO
$N$ 5368 5368 5368 5368 5368 5368
R2 0.0028 0.0027 0.0024 0.1739 0.1628 0.1455
  • Note: This is the result after PSM, matched on a 1:1 basis, with other settings not differing from the baseline regression.
A.4 Optuna Parameter Range
Table 15: Hyperparameters for XGBoost Model
Parameter Range(optuna) / Value
loss_function rmse
learning_rate [0.005, 0.2] (log scale)
max_depth [1, 3]
min_child_weight [1, 5] (log scale)
subsample [0.5, 0.8]
colsample_bytree [0.5, 0.8]
reg_lambda [1, 10] (log scale)
reg_alpha [1, 10] (log scale)
tree_method hist
device cuda
num_boost_round 1000
nfold 20
early_stopping_rounds 5
seed 42
  • Note: This table shows the range of hyperparameters Optuna searches for, automatically selecting the optimal hyperparameters for the final model configuration.